# Computing High-dimensional Confidence Sets for Arbitrary Distributions

### v1

@@ 3 Proper Learning of Euclidean Balls
Next, we estimate the location of $\widehat{c}$ in the low variance directions. Let $Y^{\prime}_{\mathrm{low}}=\Pi_{\mathrm{low}}Y^{\prime}$ denote the projection of $Y^{\prime}$ onto the low variance directions. By construction, the set $Y^{\prime}_{\mathrm{low}}$ has variance at most $\lambda_{(q+1)}(\Sigma_{Y^{\prime}})\leq\sigma^{2}$ in every direction. Since the set $Y^{*}$ contains at least a $\delta$-fraction of the points in $Y^{\prime}$, by 3.4, we have $Y^{\star}_{\mathrm{low}}=\Pi_{\mathrm{low}}Y^{\star}$ has variance at most $\sigma^{2}/\delta$ in every direction. By Lemma 3.3, the mean $\mu^{\star}_{\mathrm{low}}$ of $Y^{\star}_{\mathrm{low}}$ is close to the mean $\mu^{\prime}_{\mathrm{low}}$ of $Y^{\prime}_{\mathrm{low}}$, i.e.,
@@
Since the variance of $Y^{*}_{\mathrm{low}}$ in each direction is bounded by $\sigma^{2}/\delta$, Lemma 2.1 ensures that $\mu^{\star}_{\mathrm{low}}$ is close to most points in $Y^{\star}_{\mathrm{low}}$. That is, for a fixed choice of $\gamma$, there exists a subset $\widehat{Y}\subseteq Y^{\star}$, such that $|\widehat{Y}|\geq\left(1-\gamma\right)|Y^{\star}|$, and for every points $y\in\widehat{Y}$, we have
$\|\Pi_{\mathrm{low}}y-\mu^{\star}_{\mathrm{low}}\|\leq\|\Pi_{\mathrm{low}}y-\Pi_{\mathrm{low}}c^{*}\|+\frac{2}{\sqrt{\gamma}}\sigma.$ (9)
Combining with (8), we get that for every $y\in\widehat{Y}$,
$\|\Pi_{\mathrm{low}}y-\mu^{\prime}_{\mathrm{low}}\|$ $\leq\|\Pi_{\mathrm{low}}y-\mu^{*}_{\mathrm{low}}\|+\|\mu^{*}_{\mathrm{low}}-\mu^{\prime}_{\mathrm{low}}\|$
$\leq\|\Pi_{\mathrm{low}}y-\Pi_{\mathrm{low}}c^{*}\|+\frac{2}{\sqrt{\gamma}}\sigma+\sigma\cdot\sqrt{2(1-\delta)/\delta}$
$\leq\|\Pi_{\mathrm{low}}y-\Pi_{\mathrm{low}}c^{*}\|+\sigma\left(\frac{2}{\sqrt{\gamma}}+\sqrt{2(1-\delta)/\delta}\right).$
Now, take the center $\widehat{c}$ such that $\Pi_{\mathrm{high}}\widehat{c}=\widehat{\ell}$ and $\Pi_{\mathrm{low}}\widehat{c}=\mu^{\prime}_{\mathrm{low}}$. Then, we have that for every $y\in\widehat{Y}$, the distance from $y$ to $\widehat{c}$ is at most
$\|y-\widehat{c}\|^{2}$ $=\|\Pi_{\mathrm{low}}(y-\widehat{c})\|^{2}+\|\Pi_{\mathrm{high}}(y-\widehat{c})\|^{2}$
$=\|\Pi_{\mathrm{low}}y-\mu^{\prime}_{\mathrm{low}}\|^{2}+\|\Pi_{\mathrm{high}}y-\widehat{\ell}\|^{2}$
$\leq\left(\|\Pi_{\mathrm{low}}y-\Pi_{\mathrm{low}}c^{*}\|+\sigma\left(\frac{2}{\sqrt{\gamma}}+\sqrt{2(1-\delta)/\delta}\right)\right)^{2}+\left(\|\Pi_{\mathrm{high}}y-\Pi_{\mathrm{high}}c^{*}\|+\tau\right)^{2}$
$\leq\left(\|y-c^{*}\|+2\left(\sigma\left(\frac{2}{\sqrt{\gamma}}+\sqrt{2(1-\delta)/\delta}\right)+\tau\right)\right)^{2}.$
Since $y\in\widehat{Y}\subseteq Y^{*}$, we have
$\|y-\widehat{c}\|$ $\leq\|y-c^{*}\|+2\sigma\left(\frac{2}{\sqrt{\gamma}}+\sqrt{2(1-\delta)/\delta}\right)+2\tau$
$\leq R^{*}+2\sigma\left(\frac{2}{\sqrt{\gamma}}+\sqrt{2(1-\delta)/\delta}\right)+2\tau.$
Thus, Let $C$ be the set in $\mathbb{R}^{d}$ defined by $C=\{c~{}|~{}\Pi_{\mathrm{high}}c=\ell,\Pi_{\mathrm{low}}c=\mu^{\prime}_{\mathrm{low}},\ell\in L\}$. This list $C$ ensures that $|C|=|L|\leq\left(\frac{3R^{\prime}}{\tau}\right)^{q}$, and there exists a point $\widehat{c}\in C$ such that for the radius
$\widehat{R}=R^{\star}+2\sigma\left(\frac{2}{\sqrt{\gamma}}+\sqrt{2(1-\delta)/\delta}\right)+2\tau,$
the ball $\widehat{B}=B(\widehat{c},\widehat{R})$ contains at least $(1-\gamma)|Y^{\star}|$ points in $Y^{*}$. ∎
@@
Algorithm Dense_Ball Input: a set of point $Y\in\mathbb{R}^{d}$, target fraction $\delta\in(0,1)$, and slack $\gamma\in(0,1)$ Output: an ball $\widehat{B}\subset\mathbb{R}^{d}$ 1. Create a list of coarse balls $\mathfrak{B}=\left\{B(y_{1},\|y_{1}-y_{2}\|)~{}|~{}y_{1},y_{2}\in Y\right\}$ and remove from $\mathfrak{B}$ all balls that contain less than $\delta n$ points in $Y$. 2. Set $R_{\min}$ to be the minimum radius among all balls in $\mathfrak{B}$. Then remove from $\mathfrak{B}$ all balls with a radius greater than $2R_{\min}$. 3. Create a list of refined balls $\mathfrak{B}_{2}$ as follows: (a) Create a list of candidate centers $C_{B^{\prime}}$ for each ball $B^{\prime}$ in $\mathfrak{B}$ as follows. (b) Compute the eigenvalues $\lambda_{1},\dots,\lambda_{d}$ and eigenvectors $v_{1},\dots,v_{d}$ of the covariance $\Sigma_{Y^{\prime}}$ of points $Y^{\prime}=Y\cap B^{\prime}$. (c) Set $\Pi_{\mathrm{high}}$ be the projection onto the eigenspace of $\Sigma_{Y^{\prime}}$ with eigenvalues greater than $(R^{\prime})^{2}\log\log d/\log d$ and $\Pi_{\mathrm{low}}=I-\Pi_{\mathrm{high}}$. Set $B^{\prime}_{\mathrm{high}}=\Pi_{\mathrm{high}}B^{\prime}$ and $Y^{\prime}_{\mathrm{low}}=\Pi_{\mathrm{low}}Y^{\prime}$. Note that there are at most $q=\log d/\log\log d$ eigenvalues of $\Sigma_{Y^{\prime}}$ is greater than $(R^{\prime})^{2}\log\log d/\log d$. When $q>0$, we run a grid search over these high-variance directions; however, when $q=0$, no grid search is needed. (d) Set the list $C_{B^{\prime}}=\{c:\Pi_{\mathrm{high}}c=\ell,\Pi_{\mathrm{low}}c=\mu^{\prime}_{\mathrm{low}},\ell\in L\}$, where $\mu^{\prime}_{\mathrm{low}}$ is the mean of $Y^{\prime}_{\mathrm{low}}$ and $L$ is the $\tau$-net of the ball $B^{\prime}_{\mathrm{high}}$ for $\tau=R_{\min}/\sqrt{\log d}$. (e) Combine all candidate centers $\mathfrak{C}=\cup_{B^{\prime}\in\mathfrak{B}}C_{B^{\prime}}$. (f) Set the list $\mathfrak{B}_{2}=\left\{B(c,||y-c||)~{}|~{}c\in\mathfrak{C},y\in Y\right\}$. 4. Return the ball $\widehat{B}\in\mathfrak{B}_{2}$ with the smallest volume that contains at least $\delta(1-\gamma)$ points in $Y$.
@@
$R^{\star}\leq R_{\mathrm{min}}\leq 2\cdot R^{\star}.$
We filter $\mathfrak{B}$ to only contain balls that contain at least $\delta n$ points from $Y$, and have radius at most $2\cdot R_{\mathrm{min}}\leq 4\cdot R^{\star}$. Finding $R_{\mathrm{min}}$ and filtering the list take $\widetilde{O}(n^{3})$ time.
Now we apply the coarse-to-fine refinement of Lemma 3.6 to each ball in (the filtered) $\mathfrak{B}$, to generate a list of candidate centers for $\widehat{B}$. In particular, we apply the procedure for $q=\frac{\log d}{\log\log d}$ and $\tau=\frac{R_{\mathrm{min}}}{\sqrt{\log d}}$. For each ball $B$ in the filtered $\mathfrak{B}$, we have that the radius $R$ of $B$ is at most $2R_{\mathrm{min}}$. Thus, by Lemma 3.6, the size of the list of potential centers $C_{B}$ that we generate for $B$ is at most
$|C_{B}|\leq\left(\frac{3R^{\prime}}{\tau}\right)^{q}\leq\left(6\sqrt{\log d}\right)^{\frac{\log d}{\log\log d}}\leq d^{\frac{7}{2}}.$
@@
Now we argue that $\mathfrak{B}_{2}$ contains a ball that is a good approximation to $B^{\star}$ in volume. Consider the ball $B^{\prime}$ from (10). We have that $B^{\prime}$ contains all points in $Y^{\star}$, has radius $R^{\prime}\leq 2R^{\star}$, and is in the list $\mathfrak{B}$. Let $Y^{\prime}=Y\cap B^{\prime}$ be the subset of points of $Y$ that are captured by $B^{\prime}$. An averaging argument (Lemma 3.7) tells us that $\lambda_{(q+1)}(\Sigma_{Y^{\prime}})\leq\frac{R^{\prime 2}}{q}$, where $\Sigma_{Y^{\prime}}$ is the covariance of the points $Y^{\prime}$, and $\lambda_{(q+1)}(\Sigma_{Y^{\prime}})$ is the $(q+1)$-th largest eigenvector of $\Sigma_{Y^{\prime}}$. Thus, when we apply the procedure of Lemma 3.6 to $B^{\prime}$ with $q=\frac{\log d}{\log\log d}$ and $\tau=\frac{R_{\mathrm{min}}}{\sqrt{\log d}}$, we also have that the eigenvalue $\lambda_{(q+1)}(\Sigma_{Y^{\prime}})$ is upper bounded by $\sigma^{2}\leq\frac{R^{\prime 2}}{q}\leq\frac{(2R^{\star})^{2}\log\log d}{\log d}$. Thus, it is guaranteed that the list of potential centers $C_{B^{\prime}}$ that we get for this ball $B^{\prime}$ contains a center $\widehat{c}\in C_{B^{\prime}}$ such that $\widehat{B}=B(\widehat{c},\widehat{R})$ contains at least $(1-\gamma)|Y^{\star}|$ points for
$\widehat{R}$ $=R^{\star}+2\sigma\left(\frac{2}{\sqrt{\gamma}}+\sqrt{2(1-\delta)/\delta}\right)+2\tau$
$\leq R^{\star}+2\sqrt{\frac{(2R^{\star})^{2}\log\log d}{\log d}}\left(\frac{2}{\sqrt{\gamma}}+\sqrt{2(1-\delta)/\delta}\right)+2\frac{R_{\mathrm{min}}}{\sqrt{\log d}}$
$\leq R^{\star}+R^{\star}\sqrt{\frac{\log\log d}{\log d}}\cdot 4\left(\frac{2}{\sqrt{\gamma}}+\sqrt{2(1-\delta)/\delta}\right)+R^{\star}\cdot\frac{4}{\sqrt{\log d}}$
$\leq R^{\star}\left(1+O\left(\sqrt{\frac{\log\log d}{\log d}}\left(\frac{1}{\sqrt{\gamma}}+\sqrt{(1-\delta)/\delta}\right)\right)\right).$
@@ 3.2 Better volume approximation guarantee for almost isotropic distributions
Our strategy starts by finding a ball in $B^{\prime}$ that contains all of the points in $B^{\star}$ and has radius at most twice that of $B^{\star}$. (More precisely, we find a list of candidate balls, which can be done simply.) Trivially, this means that the standard deviation of the points in $B^{\prime}$ in every direction, is at most $R^{\star}$. Lemma 3.6 tells us that choosing the mean of the points in $B^{\prime}$ as the center of $\widehat{B}$ essentially allows us to capture most of the points in $B^{\star}$ with radius that is $R^{\star}$ plus the maximum standard deviation in any direction. This gives a $2$-approximation to the radius. We note that even a slightly stronger bound on the standard deviation can improve the approximation factor to $1+o(1)$. In Theorem 3.1, we do this by grid-searching the $\log d$ many highest variance directions of the points in $B^{\prime}$, and taking the mean in the low-variance directions. This allows us to argue about the standard deviation of the $(\log d+1)$th highest variance direction, which can be at most $\frac{R^{\star}}{\sqrt{\log d}}$ (Lemma 3.7).
This bound on the variance that we use in Theorem 3.1 is the best that our approach achieves, since it is possible that the variance of the points is concentrated in a few directions. However, if the points in $B^{\star}$ (the “inliers”) are not worst-case, and are instead approximately isotropic, we expect the standard deviation in any direction to be bounded by $\frac{R^{\star}}{\sqrt{d}}$, which we can use to give a better bound on the radius of the ball that we output.
This observation is also useful as a subroutine for our algorithm that finds a small volume ellipsoid in Section 4. At a high level, that algorithm will first estimate a linear transformation that limits the variance of the points in $B^{\star}$. Then, after applying this transformation, we can use Algorithm Dense_Ball_Isotropic to search for a ball in the transformed space with a better volume approximation.
@@
Let $\mu^{\star}$ be the mean of the points in $Y^{\star}$. We know that $Y^{\star}$ has covariance bounded by $\beta(R^{\star})^{2}/dI$, which can be estimated by $\beta(R_{\mathrm{min}})^{2}/dI$. The polynomial-time algorithm for list-decodable mean estimation (Diakonikolas et al., 2021; Diakonikolas and Kane, 2023) outputs a list $L$ of length $O(\log(1/p)/\delta)$ such that with probability $1-p$, there exists a $\widehat{\mu}\in L$ such that
$||\widehat{\mu}-\mu^{\star}||=O(\delta^{-1/2}(\beta/d)^{1/2}R_{\mathrm{min}})=O(\delta^{-1/2}(\beta/d)^{1/2}R^{\star}).$
@@
$\|y-\mu^{\star}\|\leq\|y-c^{\star}\|+2\gamma^{-1/2}(\beta/d)^{1/2})R^{\star}.$
Thus by triangle inequality, there exists an $\widehat{\mu}\in L$, such that for points $y\in\widehat{Y}$ we have
$\|y-\widehat{\mu}\|$ $\leq\|y-\mu^{\star}\|+\|\mu^{\star}-\widehat{\mu}\|$
$\leq\|y-c^{\star}\|+2\gamma^{-1/2}(\beta/d)^{1/2}R^{\star}+O(\delta^{-1/2}\beta^{1/2}R^{\star})$
$\leq\|y-c^{\star}\|+O((\gamma^{-1/2}+\delta^{-1/2})(\beta/d)^{1/2}R^{\star})$
$\leq R^{\star}\left(1+O((\gamma^{-1/2}+\delta^{-1/2})(\beta/d)^{1/2})\right).$
@@
Since $|\mathfrak{B}_{2}|\leq|L|\cdot|Y|$, we have that this list is polynomially sized. We output the minimum volume ball $\widehat{B}\in\mathfrak{B}_{2}$ such that $|\widehat{B}\cap Y|\geq(1-\gamma)\delta|Y|$. This guarantees that $\widehat{B}$ has radius $\widehat{R}$ such that
$\widehat{R}\leq R^{\star}\left(1+O((\gamma^{-1/2}+\delta^{-1/2})\frac{\beta^{1/2}}{d^{1/2}})\right).$
4 Finding Confidence Sets $(\Gamma=1+d^{-1/3+o(1)})$-Competitive with Euclidean Balls
In this section, we provide an algorithm to find a confidence set whose volume is within a $\exp(O(d^{2/3+o(1)}))$-factor of the volume of the smallest ball that contains a $\delta$ fraction of $Y$.

### v2

@@ 3 Proper Learning of Euclidean Balls
Next, we estimate the location of $\mu^{\star}$ in the low variance directions. Let $Y^{\prime}_{\mathrm{low}}=\Pi_{\mathrm{low}}Y^{\prime}$ denote the projection of $Y^{\prime}$ onto the low variance directions. By construction, the set $Y^{\prime}_{\mathrm{low}}$ has variance at most $\lambda_{(q+1)}(\Sigma_{Y^{\prime}})\leq\sigma^{2}$ in every direction. Since the set $Y^{*}$ contains at least a $\delta$-fraction of the points in $Y^{\prime}$, by 3.4, we have $Y^{\star}_{\mathrm{low}}=\Pi_{\mathrm{low}}Y^{\star}$ has variance at most $\sigma^{2}/\delta$ in every direction. By Lemma 3.3, the mean $\mu^{\star}_{\mathrm{low}}$ of $Y^{\star}_{\mathrm{low}}$ is close to the mean $\mu^{\prime}_{\mathrm{low}}$ of $Y^{\prime}_{\mathrm{low}}$, i.e.,
@@
Since the variance of $Y^{\star}_{\mathrm{low}}$ is bounded by $\sigma^{2}/\delta$, and $\mu^{\prime}_{\mathrm{low}}$ is close to $\mu^{\star}_{\mathrm{low}}$, Lemma 2.1 tells us that, for a fixed choice of $\gamma$, there exists a subset $\widehat{Y}\subseteq Y^{\star}$ such that $|\widehat{Y}|\geq(1-\gamma)|Y^{\star}|$, and for every point $y\in\widehat{Y}$,
$||\Pi_{\mathrm{low}}y-\mu^{\prime}_{\mathrm{low}}||^{2}_{2}\leq\|\Pi_{\mathrm{low}}(y-c^{\star})\|_{2}^{2}+\frac{\sigma^{2}}{\gamma}\left(1+\frac{2(1-\delta)}{\delta}\right).$
For $q=0$, this gives the final bound.
For $q>0$, we take $\widehat{c}$ such that $\Pi_{\mathrm{high}}\widehat{c}=\widehat{\ell}$ and $\Pi_{\mathrm{low}}\widehat{c}=\mu^{\prime}_{\mathrm{low}}$. Putting the bounds together, we have that there exists a set $\widehat{Y}\subseteq Y^{\star}$, with $|\widehat{Y}|\geq(1-\gamma)|Y^{\star}|$, such that for every $y\in\widehat{Y}$,
$\|y-\widehat{c}||_{2}^{2}$ $=||\Pi_{\mathrm{low}}(y-\widehat{c})||_{2}^{2}+||\Pi_{\mathrm{high}}(y-\widehat{c})||_{2}^{2}$
$\leq||\Pi_{\mathrm{low}}(y-c^{\star})||_{2}^{2}+\frac{\sigma^{2}}{\gamma}\left(1+\frac{2(1-\delta)}{\delta}\right)+\left(\|\Pi_{\mathrm{high}}(y-c^{\star})\|_{2}+\tau\right)^{2}$
$\leq\left(||y-c^{\star}||_{2}+O\left(\frac{\sigma}{\sqrt{\gamma\delta}}\right)+\tau\right)^{2},$
where both the tighter bound (second line) and looser bound (third line) will be useful in applications of the lemma.
@@
Algorithm Dense_Ball Input: a set of point $Y\in\mathbb{R}^{d}$, target fraction $\delta\in(0,1)$, and slack $\gamma\in(0,1)$ Output: an ball $\widehat{B}\subset\mathbb{R}^{d}$ 1. Create a list of coarse balls $\mathfrak{B}=\left\{B(y_{1},\|y_{1}-y_{2}\|)~{}|~{}y_{1},y_{2}\in Y\right\}$ and remove from $\mathfrak{B}$ all balls that contain less than $\delta n$ points in $Y$. 2. Set $R_{\min}$ to be the minimum radius among all balls in $\mathfrak{B}$. Then remove from $\mathfrak{B}$ all balls with a radius greater than $2R_{\min}$. 3. Create a list of refined balls $\mathfrak{B}_{2}$ as follows: (a) Create a list of candidate centers $C_{B^{\prime}}$ for each ball $B^{\prime}$ in $\mathfrak{B}$ as follows. (b) Compute the eigenvalues $\lambda_{1},\dots,\lambda_{d}$ and eigenvectors $v_{1},\dots,v_{d}$ of the covariance $\Sigma_{Y^{\prime}}$ of points $Y^{\prime}=Y\cap B^{\prime}$. (c) Set $\Pi_{\mathrm{high}}$ be the projection onto the eigenspace of $\Sigma_{Y^{\prime}}$ with eigenvalues greater than $(R^{\prime})^{2}\log\log d/\log d$ and $\Pi_{\mathrm{low}}=I-\Pi_{\mathrm{high}}$. Set $B^{\prime}_{\mathrm{high}}=\Pi_{\mathrm{high}}B^{\prime}$ and $Y^{\prime}_{\mathrm{low}}=\Pi_{\mathrm{low}}Y^{\prime}$. Note that there are at most $q=\log d/\log\log d$ eigenvalues of $\Sigma_{Y^{\prime}}$ is greater than $(R^{\prime})^{2}\log\log d/\log d$. When $q>0$, we run a grid search over these high-variance directions; however, when $q=0$, no grid search is needed. (d) Set the list $C_{B^{\prime}}=\{c:\Pi_{\mathrm{high}}c=\ell,\Pi_{\mathrm{low}}c=\mu^{\prime}_{\mathrm{low}},\ell\in L\}$, where $\mu^{\prime}_{\mathrm{low}}$ is the mean of $Y^{\prime}_{\mathrm{low}}$ and $L$ is the $\tau$-net of the ball $B^{\prime}_{\mathrm{high}}$ for $\tau=R_{\min}/\log d$. (e) Combine all candidate centers $\mathfrak{C}=\cup_{B^{\prime}\in\mathfrak{B}}C_{B^{\prime}}$. (f) Set the list $\mathfrak{B}_{2}=\left\{B(c,||y-c||)~{}|~{}c\in\mathfrak{C},y\in Y\right\}$. 4. Return the ball $\widehat{B}\in\mathfrak{B}_{2}$ with the smallest volume that contains at least $\delta(1-\gamma)$ points in $Y$.
@@
$R^{\star}\leq R_{\mathrm{min}}\leq 2R^{\star}.$
We filter $\mathfrak{B}$ to only contain balls that contain at least $\delta n$ points from $Y$, and have radius at most $2R_{\mathrm{min}}\leq 4R^{\star}$. Finding $R_{\mathrm{min}}$ and filtering the list take $\widetilde{O}(n^{3})$ time.
Now we apply the coarse-to-fine refinement of Lemma 3.6 to each ball in (the filtered) $\mathfrak{B}$, to generate a list of candidate centers for $\widehat{B}$. In particular, we apply the procedure for $q=\frac{\log d}{\log\log d}$ and $\tau=\frac{R_{\mathrm{min}}}{\log d}$. For each ball $B$ in the filtered $\mathfrak{B}$, we have that the radius $R$ of $B$ is at most $2R_{\mathrm{min}}$. Thus, by Lemma 3.6, the size of the list of potential centers $C_{B}$ that we generate for $B$ is at most
$|C_{B}|\leq\left(\frac{3R^{\prime}}{\tau}\right)^{q}\leq\left(6\log d\right)^{\frac{\log d}{\log\log d}}\leq d^{3}.$
@@
Now we argue that $\mathfrak{B}_{2}$ contains a ball that is a good approximation to $B^{\star}$ in volume. Consider the ball $B^{\prime}$ from (10). We have that $B^{\prime}$ contains all points in $Y^{\star}$, has radius $R^{\prime}\leq 2R^{\star}$, and is in the list $\mathfrak{B}$. Let $Y^{\prime}=Y\cap B^{\prime}$ be the subset of points of $Y$ that are captured by $B^{\prime}$. An averaging argument (Lemma 3.7) tells us that $\lambda_{(q+1)}(\Sigma_{Y^{\prime}})\leq\frac{R^{\prime 2}}{q}$, where $\Sigma_{Y^{\prime}}$ is the covariance of the points $Y^{\prime}$, and $\lambda_{(q+1)}(\Sigma_{Y^{\prime}})$ is the $(q+1)$-th largest eigenvector of $\Sigma_{Y^{\prime}}$. Thus, when we apply the procedure of Lemma 3.6 to $B^{\prime}$ with $q=\frac{\log d}{\log\log d}$ and $\tau=\frac{R_{\mathrm{min}}}{\log d}$, we also have that the eigenvalue $\lambda_{(q+1)}(\Sigma_{Y^{\prime}})$ is upper bounded by $\sigma^{2}\leq\frac{R^{\prime 2}}{q}\leq\frac{(2R^{\star})^{2}\log\log d}{\log d}$. This guarantees that, for the list of potential centers $C_{B^{\prime}}$ that we get for this ball $B^{\prime}$, there is a center $\widehat{c}\in C_{B^{\prime}}$ and a subset $\widehat{Y}\subseteq Y^{\star}$, $|\widehat{Y}|\geq(1-\gamma)|Y^{\star}|$, such that for every $y\in\widehat{Y}$,
$\|y-c^{\star}\|_{2}^{2}$ $\leq\|\Pi_{\mathrm{low}}(y-c^{\star})\|_{2}^{2}+O\Big{(}\frac{\sigma^{2}}{\gamma\delta}\Big{)}+(\|\Pi_{\mathrm{high}}(y-c^{\star})\|_{2}+\tau)^{2},$
where $\Pi_{\mathrm{high}}$ is the projection onto the top-$q$ eigenspace of $\Sigma_{Y^{\prime}}$ and $\Pi_{\mathrm{low}}$ is the projection onto the space orthogonal to the top-$q$ eigenspace of $\Sigma_{Y^{\prime}}$.
To bound the second term, we use the fact that
$(a+b)^{2}\leq(1+\varepsilon)a^{2}+(1+\frac{1}{\varepsilon})b^{2},~{}\forall\varepsilon>0.$
Using $\tau=\frac{R_{\mathrm{min}}}{\log d}$, and $\varepsilon=\frac{1}{\log d}$, we get
$\|y-c^{\star}\|_{2}^{2}$ $\leq\|\Pi_{\mathrm{low}}(y-c^{\star})\|_{2}^{2}+O\left(\frac{R^{\prime 2}\log\log d}{\gamma\delta\log d}\right)+\left(1+\frac{1}{\log d}\right)\|\Pi_{\mathrm{high}}(y-c^{\star})\|_{2}^{2}+(1+\log d)\cdot\frac{R_{\mathrm{min}}^{2}}{\log^{2}d}$
$\leq\left(1+\frac{1}{\log d}\right)\|y-c^{\star}\|_{2}^{2}+O\left(\frac{(R^{\star})^{2}\log\log d}{\gamma\delta\log d}\right)$
$\leq(R^{\star})^{2}\left(1+O\left(\frac{\log\log d}{\gamma\delta\log d}\right)\right).$
Thus, it is guaranteed that the list of potential centers $C_{B^{\prime}}$ that we get for this ball $B^{\prime}$ contains a center $\widehat{c}\in C_{B^{\prime}}$ such that $\widehat{B}=B(\widehat{c},\widehat{R})$ contains at least $(1-\gamma)|Y^{\star}|$ points for
$\widehat{R}$ $\leq R^{\star}\left(1+O\left(\frac{\log\log d}{\gamma\delta\log d}\right)\right)^{1/2},$
which corresponds to a volume bound of
$\operatorname*{\mathrm{vol}}(\widehat{B})$ $\leq\operatorname*{\mathrm{vol}}(B^{\star})\exp\left(O\left(\frac{d\log\log d}{\gamma\delta\log d}\right)\right).$
@@ 3.2 Better volume approximation guarantee for almost isotropic distributions
Our strategy starts by finding a ball in $B^{\prime}$ that contains all of the points in $B^{\star}$ and has radius at most twice that of $B^{\star}$. (More precisely, we find a list of candidate balls, which can be done simply.) Trivially, this means that in every direction the standard deviation of the points in $B^{\prime}$ is at most $R^{\star}$. Lemma 3.6 tells us that choosing the mean of the points in $B^{\prime}$ as the center of $\widehat{B}$ essentially allows us to capture most of the points in $B^{\star}$ with radius that is $R^{\star}$ plus the maximum standard deviation in any direction. This gives a $2$-approximation to the radius. We note that even a slightly stronger bound on the standard deviation can improve the approximation factor to $1+o(1)$. In Theorem 3.1, we do this by grid-searching the $\log d$ many highest variance directions of the points in $B^{\prime}$, and taking the mean in the low-variance directions. This allows us to argue about the standard deviation of the $(\log d+1)$th highest variance direction, which can be at most $\frac{R^{\star}}{\sqrt{\log d}}$ (Lemma 3.7).
This bound on the variance that we use in Theorem 3.1 is the best that our approach achieves, since it is possible that the variance of the points is concentrated in a few directions. However, if the points in $B^{\star}$ (the “inliers”) are not worst-case, and are instead approximately isotropic, we expect the standard deviation in any direction to be bounded by $\frac{R^{\star}}{\sqrt{d}}$, which we can use to give a better bound on the radius of the ball that we output. In our earlier argument, we bounded the variance not only of the inliers– the points in $B^{\star}$– but also all of $Y^{\prime}$, which includes some of the outliers. To utilize this weaker assumption on the variance, we appeal to algorithms for list-decodable robust mean estimation, which work exactly in this setting where the inliers are well-behaved, but the outliers may be arbitrary.
This observation is also useful as a subroutine for our algorithm that finds a small volume ellipsoid in Section 4. At a high level, that algorithm will first estimate a linear transformation that limits the variance of the points in $B^{\star}$. Then, after applying this transformation, we essentially use Algorithm Dense_Ball_Isotropic to search for a ball in the transformed space with a better volume approximation. In that application, we will apply the transformation to and control the variance of all of our points, inliers and outliers, and thus avoid the need for list-decodable robust mean estimation.
@@
Let $\mu^{\star}$ be the mean of the points in $Y^{\star}$. We know that $Y^{\star}$ has covariance p.s.d. dominated by $\tfrac{\beta(R^{\star})^{2}}{d}I$, which is in turn p.s.d. dominated by $\tfrac{\beta(R_{\mathrm{min}})^{2}}{d}I$. The polynomial-time algorithm for list-decodable mean estimation (Diakonikolas et al., 2021; Diakonikolas and Kane, 2023) outputs a list $L$ of length $O(\log(1/p)/\delta)$ such that with probability $1-p$, there exists a $\widehat{\mu}\in L$ such that
$||\widehat{\mu}-\mu^{\star}||=O\Big{(}\Big{(}\frac{\beta}{\delta d}\Big{)}^{1/2}R_{\mathrm{min}}\Big{)}=O\Big{(}\Big{(}\frac{\beta}{\delta d}\Big{)}^{1/2}R^{\star}\Big{)}.$
@@
$\|y-\widehat{\mu}\|^{2}_{2}$ $\leq\|y-c^{\star}\|_{2}^{2}+\gamma^{-1}\left(\frac{\beta}{d}(R_{\mathrm{min}})^{2}+O\left(\delta^{-1}(\beta/d)(R^{\star})^{2}\right)\right)$
$\leq(R^{\star})^{2}+O\left(\gamma^{-1}\delta^{-1}(\beta/d)(R^{\star})^{2}\right)$
$\leq(R^{\star})^{2}\left(1+O\left(\frac{\beta}{\gamma\delta d}\right)\right)$
@@
Since $|\mathfrak{B}_{2}|\leq|L|\cdot|Y|$, we have that this list is polynomially sized. We output the minimum volume ball $\widehat{B}\in\mathfrak{B}_{2}$ such that $|\widehat{B}\cap Y|\geq(1-\gamma)\delta|Y|$. This guarantees that $\widehat{B}$ has radius $\widehat{R}$ with
$\widehat{R}\leq R^{\star}\sqrt{1+O\left(\frac{\beta}{\gamma\delta d}\right)}.$
4 Finding Confidence Sets $(\Gamma=1+d^{-1/2+o(1)})$-Competitive with Euclidean Balls
In this section, we provide an algorithm to find a confidence set whose volume is within a
$\exp(O(d^{1/2+o(1)}))$-factor of the volume of the smallest ball that contains a $\delta$ fraction of $Y$.